OpenStudy (anonymous):

dy/dx=-3y-2; y(2) = 0 ... when solved the equation i have got y= -2/3 +Ke^-3t now what shall i do with y(2) = 0...pls help me

hartnn (hartnn):

you plug in x=2 in the equation you got to find the value of k

hartnn (hartnn):

and put y=0

OpenStudy (anonymous):

i didnt understand , you mean in the question pls explain

hartnn (hartnn):

not in the question, the equation u got y= -2/3 +Ke^-3t y(2) = 0 means , when t=2, y=0 so plug in t=2,y=0 in y= -2/3 +Ke^-3t

hartnn (hartnn):

your independent variable is "x" right ? so you should have got y= -2/3 +Ke^-3x isn't it ?? hence y(2) = 0 means , when x=2, y=0 so plug in x=2,y=0 in y= -2/3 +Ke^-3x

hartnn (hartnn):

got it ?

OpenStudy (anonymous):

yes.. thanks... let me try

hartnn (hartnn):

k, take your time :)

OpenStudy (anonymous):

now 0=-2/3+Ke^-6 am i right..

hartnn (hartnn):

yes, try to isolate k now

OpenStudy (anonymous):

ok... i am trying...:)

OpenStudy (unklerhaukus):

that's right you almost there,

OpenStudy (unklerhaukus):

when you get K, just plug it into y(x) = -2/3 +Ke^(-3x) and simplify a little bit, and that's it.

OpenStudy (anonymous):

i lost it .. i tried k=-2/3+e^-6 if its right ..then got stuck again

hartnn (hartnn):

you still had to isolate k from -2/3+ke^-6=0 start by adding 2/3 on both sides

OpenStudy (unklerhaukus):

you had 0=-2/3+Ke^-6 add 2/3 to both sides and multiply by e^6

OpenStudy (anonymous):

i am not reaching anywhere.... pls show me this then i will get an idea

hartnn (hartnn):

0=-2/3+Ke^-6 2/3 = ke^{-6} 2/3 * e^6 = k e^{-6} e^6 so, k= 2/3e^6

OpenStudy (anonymous):

in the third line u used exponential rule.. like m^n* m^m = m^n+m then that became 1 am i right ...

hartnn (hartnn):

yes, exactly -6+6=0 e^0=1 i multiplied by e^6 to make the exponent =0

OpenStudy (anonymous):

ok now its clear for me.. wow my brain so relaxed.. a million thanks will come back soon with an another question :)

hartnn (hartnn):

you are most welcome ^_^ oh, and bdw \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]